OpenStudy (anonymous):

y varies inversely as twice x. When x = 1, y = 3. Find y when x = 4.

OpenStudy (anonymous):

Answer Choices a.) one-half b.) two-thirds c.) three-fourths d.) one-fourth

OpenStudy (anonymous):

y = k/x nd 3 = k/1 find k nd then y = k/4

OpenStudy (anonymous):

y varies inversely as twice x: y ~ \(\large \frac{k}{2x}\rightarrow 3=\frac{k}{2(1)}\rightarrow k=6\) so \(\large y=\frac{6}{2x} \) now just replace x with 4 and figure out y.....

OpenStudy (anonymous):

can someone please give me the answer :( please i dont understand

OpenStudy (ganpat):

y = k/2x 1. for x =1 & y =3 3 = k / 2 (1) , so K = 6 2. Now x = 4, y = ?? y = 6/ 2(x) y = 6/ 2 (4) y = 3/4... is the answer..

OpenStudy (anonymous):

THANK YOU SO MUCH! it makes much sense now.

OpenStudy (ganpat):

wc.. :)